Two squares have sides of lengths x cm and( 2x + l) cm. The sum of the perimeters of the two squares equals 88 cm. The length of each side of the larger square is
(a) 7 (b) 15 (c) 29 (d) 57 (e) none of these

Answer by stanbon(75887)

Two squares have sides of lengths x cm and( 2x + l) cm. The sum of the perimeters of the two squares equals 88 cm. The length of each side of the larger square is
------------
1st perimeter = 4x cm
2nd perimeter = 4(2x+1) = 8x+4 cm
---------------------
Equation:
4x + 8x+4 = 88
12x = 84
x = 7
-----
1st side: x = 7
2nd side: 2x+1 = 15
=====================
Cheers,
Stan H.






(a) 7 (b) 15 (c) 29 (d) 57 (e) none of these